# CausAdv: A Causal-based Framework for Detecting Adversarial Examples

### v2

@@ Abstract
Deep learning has led to tremendous success in computer vision, largely due to Convolutional Neural Networks (CNNs). However, CNNs have been shown to be vulnerable to crafted adversarial perturbations. This vulnerability of adversarial examples has has motivated research into improving model robustness through adversarial detection and defense methods. In this paper, we address the adversarial robustness of CNNs through causal reasoning. We propose CausAdv: a causal framework for detecting adversarial examples based on counterfactual reasoning. CausAdv learns both causal and non-causal features of every input, and quantifies the counterfactual information (CI) of every filter of the last convolutional layer. We then perform a statistical analysis of the filters’ CI across clean and adversarial samples, to demonstrate that adversarial examples exhibit different CI distributions compared to clean samples. Our results show that causal reasoning enhances the process of adversarial detection without the need to train a separate detector. Moreover, we illustrate the efficiency of causal explanations as a helpful detection tool by visualizing the extracted causal features.
@@
organization=Laboratory of Informatics and applications of M’sila, Faculty of Mathematics and Computer Science, University of M’sila,addressline=PO Box 166, city=Ichebilia, postcode=28000, state=M’sila, country=Algeria
@@ 1 Introduction
The act of generating adversarial examples to threat the security of DNNs systems is called adversarial attacks. Adversarial attacks range in different categories. Regarding adversary’s knowledge, we have white and black box attacks. For white box attacks, the adversary has a full knowledge about the trained classification model including model architecture, hyperparameters, activation functions, and model weights. These attacks usually take the gradient of the predictions probabilities with respect to particular pixels. Some famous white box attacks are FGSM [19], PGD [20], BIM [21], Square[60] and C&W [22].
@@ 2 Related work
Although adversarial training (AT) remains the most effective strategy, the internal activation behavior of DNNs under small perturbations is still not well understood, particularly how minor changes accumulate across layers to fool the network and how to be mitigated. To address this, several studies have proposed robust architectural designs [24, 25, 26]. One of the most used methods is gradient masking, which aims to reduce sensitivity of DNNs to small perturbations. Papernot et al. [24] introduced the defensive distillation methods, which involves training a large model, called the teacher, to maximize the likelihood of correct class predictions, and then transfer this knowledge to a smaller model, called the student. The student model is then expected to predict the same probability distribution over classes. So, relying on a hard softmax function could provide smooth decision boundaries, which makes it very hard to find adversarial examples without losing imperceptibility. While effective against attacks such as JSMA [38] on MNIST and CIFAR, this approach remains vulnerable to transferable black-box attacks. Gu and Rigazio [25] proposed gradient regularization through contractive autoencoder trained on adversarial examples prior to classification. This method imposes a layer-wise contractive penalty to constrain the model’s sensitivity. However, the effectiveness of this approach depends on the class of adversarial examples used during training, and in some cases, the combined autoencoder-classifier model may fail.
@@ 4.1 Experiment Setup
We evaluate CausAdv on both ImageNet and CIFAR-10 against both targeted and untargeted variants of four widely used $l\infty$ bounded adversarial attacks: Fast Gradient Sign Method (FGSM)[19], Projected Gradient Descent (PGD)[20], Basic Iterative Method (BIM)[21] and Square Attack [60]. While the first tree methods can be configured to be both targeted and untargeted, Suqare attack is originally untargeted. For ImageNet targeted attacks, we randomly assign unique target labels, one for each of the 1k selected source classes, ensuring that no target label overlaps with its corresponding source class. A similar procedure is used for CIFAR-10, where each of the 10,000 samples is assigned a distinct target class (different from its true label), covering all 10 classes.
To ensure the attacks truly change the model’s prediction, we carefully select the perturbation budget $\epsilon$. For ImageNet, we found that $\epsilon=4$ was insufficient to reliably induce misclassification across all attacks. Therefore, we use $\epsilon=8$, which achieved an acceptable attack success rate (ASR) for both targeted and untargeted variants across all attacks.
@@ 4.2 Experimental results
As shown in the results, $4,211$ of $6,000$ clean samples have their true class among the top-5 most correlated prototypes, which proves the usefulness of the Pearson correlation measure over CI distributions. In contrast, adversarial samples generally exhibit lower top-5 correlation counts except in the case of the Square attack, which yields the highest number of matching samples at $2041$. However, it is worth noting that given the perturbation budget $\epsilon=8$ Square attack does not attend a high ASR comparing to other methods, and this justifies the high number of top-5 correlated labels. To ensure a fair comparison, it is important to note that the $6,000$ samples were randomly selected.
@@
4.3 Comparison with some detection methods
Table 5 presents the results of CausAdv compared to four defense methods: Feature Squeezing (FS)[33], Spatial Smoothing (SS) [33], Gaussian Data Augmentation (GDA) [64] and CutMix [65], evaluated on the same set of $6,000$ adversarial samples from the ImageNet dataset. We employ the True Positive Rate (TPR) or recall as a measure of performance on the adversarial examples. All defense methods were implemented using the Adversarial Robustness Toolbox (ART) library [62], under consistent experimental settings. As shown in the results, for a perturbation budget of $\epsilon=8$, most defense methods fail to recover the correct classes for the adversarial samples, whether the attacks are targeted or untargeted, except for Square attack, which has the highest number of samples being recovered with GDA. This is due mainly to the low ASR of this attack in the first place compared to other attacks as reported in Table  1. By comparing defense methods, Spatial Smoothing (SS) performs slightly better than the other methods, retrieving a higher number of correct predictions, specifically for PGD and BIM attacks but its effectiveness remains limited overall.
These findings highlight the difficulty of defending against adversarial examples on large-scale datasets like ImageNet. In contrast, our proposed method, CausAdv, significantly outperforms all tested defenses, particularly in handling BIM and FGSM-targeted attacks. All experiments were conducted under identical conditions using ART to ensure a fair comparison.
Attack Type Attack Defense Method
@@ 4.3 Comparison with some detection methods
Targeted FGSM 504 515 515 517 91.43%
PGD 729 1,880 454 442 85.75%
BIM 670 1,813 409 411 100%
Untargeted FGSM 404 654 385 387 81.68
PGD 1,012 1,630 1,012 629 82.80%
BIM 161 262 155 155 100%
Square 2,461 2,069 2,593 2,544 42.40%
Table 5: True Positive Rate(TPR) of each defense method for both targeted and untargeted attacks ($\epsilon=8$) given 6,000 samples.

### v3

@@ Abstract
Deep learning has led to tremendous success in computer vision, largely due to Convolutional Neural Networks (CNNs). However, CNNs have been shown to be vulnerable to crafted adversarial perturbations. This vulnerability of adversarial examples has has motivated research into improving model robustness through adversarial detection and defense methods. In this paper, we address the adversarial robustness of CNNs through causal reasoning. We propose CausAdv: a causal framework for detecting adversarial examples based on counterfactual reasoning. CausAdv learns both causal and non-causal features of every input, and quantifies the counterfactual information (CI) of every filter of the last convolutional layer. We then perform a statistical analysis of the filters’ CI across clean and adversarial samples, to demonstrate that adversarial examples exhibit different CI distributions compared to clean samples. Our results show that causal reasoning enhances the process of adversarial detection without the need to train a separate detector. Moreover, we illustrate the efficiency of causal explanations as a helpful detection tool by visualizing the extracted causal features. Code for reproducing our results is available at : https://github.com/HichemDebbi/CausAdv/tree/main.
@@
organization=Laboratory of Informatics and applications of M’sila, Department of Computer Science, University of M’sila,addressline=PO Box 166, city=Ichebilia, postcode=28000, state=M’sila, country=Algeria
@@ 1 Introduction
The act of generating adversarial examples to threat the security of DNNs systems is called adversarial attacks. Adversarial attacks range in different categories. Regarding adversary’s knowledge, we have white and black box attacks. For white box attacks, the adversary has a full knowledge about the trained classification model including model architecture, hyperparameters, activation functions, and model weights. These attacks usually take the gradient of the predictions probabilities with respect to particular pixels. Some famous white box attacks are FGSM [22], PGD [37], BIM [32], Square[2] and C&W [8]. While most adversarial attacks are white-box attacks, adversarial examples can nevertheless be transferred to black-box settings due to their transferability property[42]. Several works have further improved this transferability by proposing more sophisticated attack strategies [64].
@@ 2 Related work
Although adversarial training (AT) remains the most effective strategy [45], the internal activation behavior of DNNs under small perturbations is still not well understood, particularly how minor changes accumulate across layers to fool the network and how to be mitigated. To address this, several studies have proposed robust architectural designs [43, 24, 62]. One of the most used methods is gradient masking, which aims to reduce sensitivity of DNNs to small perturbations. Papernot et al. [43] introduced the defensive distillation methods, which involves training a large model, called the teacher, to maximize the likelihood of correct class predictions, and then transfer this knowledge to a smaller model, called the student. The student model is then expected to predict the same probability distribution over classes. So, relying on a hard softmax function could provide smooth decision boundaries, which makes it very hard to find adversarial examples without losing imperceptibility. While effective against attacks such as JSMA [44] on MNIST and CIFAR, this approach remains vulnerable to transferable black-box attacks. Gu and Rigazio [24] proposed gradient regularization through contractive autoencoder trained on adversarial examples prior to classification. This method imposes a layer-wise contractive penalty to constrain the model’s sensitivity. However, the effectiveness of this approach depends on the class of adversarial examples used during training, and in some cases, the combined autoencoder-classifier model may fail.
@@ 4.1 Experiment Setup
We evaluate CausAdv on both ImageNet and CIFAR-10 against both targeted and untargeted variants of four widely used $l\infty$ bounded adversarial attacks: Fast Gradient Sign Method (FGSM)[22], Projected Gradient Descent (PGD)[37], Basic Iterative Method (BIM)[32], C&W [8] and Square Attack [2]. While the first four methods can be configured to be both targeted and untargeted, Suqare attack is originally a recent untargeted attack. For ImageNet targeted attacks, we randomly assign unique target labels, one for each of the 1k selected source classes, ensuring that no target label overlaps with its corresponding source class. A similar procedure is used for CIFAR-10, where each of the 10,000 samples is assigned a distinct target class (different from its true label), covering all 10 classes.
To ensure the attacks truly change the model’s prediction, we carefully select the perturbation budget $\epsilon$. For ImageNet, we found that $\epsilon=4$ was insufficient to reliably induce misclassification across all attacks. Therefore, we use $\epsilon=8$, which achieved an acceptable attack success rate (ASR) for both targeted and untargeted variants across all attacks. While FGSM, PGD, BIM and Square attack share a similar configuration based on the perturbation budget $\epsilon$, C&W differs in that it is an optimization-based attack and is therefore configured via the number of optimization iterations, which is set to 500. Under these settings, FGSM, PGD, and BIM achieve high ASR, whereas Square Attack and C&W attain lower ASR.
@@ 4.2 Experimental results
As shown in the results, $4,211$ of $6,000$ clean samples have their true class among the top-5 most correlated prototypes, which proves the usefulness of the Pearson correlation measure over CI distributions. In contrast, adversarial samples generally exhibit lower top-5 correlation counts except in the case of C&W and Square attack, which yield the highest number of matching samples at $2032$ and $2041$ respectively. However, it is worth noting that given the perturbation budget $\epsilon=8$, C&W and Square attack do not attend a high ASR comparing to other methods, and this justifies the high number of top-5 correlated labels. To ensure a fair comparison, it is important to note that the $6,000$ samples were randomly selected.
@@
4.3 Comparison with existing detection methods
Table 5 presents the results of CausAdv compared to four defense methods: Feature Squeezing (FS)[63], Spatial Smoothing (SS) [63], Gaussian Data Augmentation (GDA) [67] and CutMix [66], evaluated on the same set of $6,000$ adversarial samples from the ImageNet dataset. We employ the True Positive Rate (TPR) or Recall as a measure of performance on the adversarial examples. All defense methods were implemented using the Adversarial Robustness Toolbox (ART) library [3], under consistent experimental settings. As shown in the results, for a perturbation budget of $\epsilon=8$, most defense methods fail to recover the correct classes for the adversarial samples, whether the attacks are targeted or untargeted, especially for Square and C&W, which is due mainly to the low ASR of these attacks in the first place compared to other attacks as reported in Table  1, in addition to their efficiency. By comparing defense methods, Spatial Smoothing (SS) performs slightly better than the other methods, retrieving a higher number of correct predictions, specifically for PGD and BIM attacks but its effectiveness remains limited overall.
These findings highlight the difficulty of defending against adversarial examples on large-scale datasets like ImageNet. In contrast, our proposed method, CausAdv, significantly outperforms all tested defenses, particularly in handling BIM and FGSM-targeted attacks. CausAdv achieves a lower detection rate only for the Square and C&W attacks because these attacks exhibit a lower ASR compared to other attacks. Consequently, the perturbations introduced by Square and C&W are relatively weak, causing the adversarial samples to appear similar to clean samples. As a result, the causal features are not significantly altered, making them harder for CausAdv to distinguish from normal samples.
Attack Type Attack ASR (%) TPR (%) by Defense Method
@@ 4.3 Comparison with existing detection methods
Targeted FGSM 91.38 8.40 8.58 8.58 8.61 91.43
PGD 92.63 12.15 31.33 7.56 7.36 85.75
BIM 93.15 11.16 30.21 6.81 6.85 100
C&W 33.20 10.76 12.51 3.03 0.0 33.86
Untargeted FGSM 93.55 6.73 10.90 6.41 6.45 81.68
PGD 89.52 16.86 27.16 16.86 10.48 82.80
BIM 96.40 2.68 4.38 2.58 2.58 100
C&W 36.23 7.45 9.01 3.08 0.0 31.70
Square 57.52 10.11 15.73 4.34 0.0 42.44
Table 5: True Positive Rate(TPR) of each defense method for both targeted and untargeted attacks ($\epsilon=8$) given 6,000 imagenet samples.